The height of the helicopter from the ground is given by h = 3.00 t^3. Here, the unit of h is m and the unit of t is s. After t = 2.00s, a small postal carriage is dropped from the helicopter. How long does it take for the postal mail to fall and reach the ground?
solution is 7.96s why?
Answer #1

Height ha 3+3 at t= 25cc, ha 3x a3 = 94 m - Ground Velocity of helicopter VCH) a dvema (3+3) VC+) = 3 x 3+2 = 9t2 vet) = us a